A parallel combination of a 1.27 uF capacitor and a 2.67 uF one is connected in series to a 4.85 uF capacitor. This three capacitor combination is connected to a 18.9 V battery. FInd the charge of each capacitor.


A)   Charge of 4.85 uF capacitor? _______C


B)   Charge of 1.27 uF capacitor? _______C


C)   Charge of 2.67 uF capacitor? _______C
Answer #1

A)

Ceq = 1/(1/4.85 + 1/(1.27+2.67)) = 2.1739 uF

qeq = Ceq V = 2.1739 * 18.9 = 41.087 uC

q(4.85) = 41.1 uC


B)

V' = 18.9 - (41.087/4.85) = 10.428 V

q(1.27) = 1.27 * 10.428 = 13.2 uC


C)

q(2.67) = 2.67 * 10.428 =27.8 uC